# Iwahori Matsumoto presentation for modules of Iwahori fixed functions on symmetric spaces

### v2

@@ Abstract
We study the space $S(X)^{I}$ of smooth functions on a symmetric space $X=G/H$ invariant to the action of an Iwahori subgroup $I$, as a module over $\mathcal{H}(G,I)$, the Iwahori Hecke algebra of a p-adic group $G$. We present a description of this module that generalizes the description given to $\mathcal{H}(G,I)$ by Iwahori and Matsumoto.
@@ 1 Introduction
Let $F$ be a non archimedean local field, of residual characteristic not 2.
Let $\mathbf{G}$ be a connected reductive group over $F$ and let $G=\mathbf{G}(F)$. Let $\sigma:\mathbf{G}\rightarrow\mathbf{G}$ be an algebraic involution defined over $F$. We denote by $\mathbf{H}$ the fixed points of $\sigma$, and by $H$ the $F$-points of $\mathbf{H}$. We call $X=G/H$ a p-adic symmetric space.
@@
Let $I$ be an Iwahori subgroup of $G$ (defined in [IM65]). The group $I$ splits the category $\mathcal{M}(G)$ of smooth representations of $G$ (see [Ber84]). The subcategory of the representations that are generated by their $I$ fixed vectors is the so-called principal block. The category of admissible representations generated by their $I$ fixed vectors was studied in [Bor76] and was shown to be equivalent to the category of finite dimensional $\mathcal{H}(G,I)$ modules, the equivalence is given by $V\rightarrow V^{I}$. Here, $\mathcal{H}(G,I)=S(G)^{I\times I}$ is the Hecke Iwahori algebra of $G$ consisting of compactly supported locally constant $I$ bi-invariant functions on $G$.
A first step towards the study of smooth $H$-distinguished representations is the study of $\mathcal{H}(G,I)$ modules with a non-trivial equivariant map from $S(X)^{I}$. Thus a description of $S(X)^{I}$ as a module over $\mathcal{H}(G,I)$ is a crucial step. In the present work, we do exactly this. Below we describe our main results in details: Using the affine building $\mathcal{B}$ of $G$ we are able to describe the $I$ orbits on $X$ giving a linear basis of $S(X)^{I}$ over $\mathbb{C}$. Then we describe the action of the Iwahori Matsumoto generators of $\mathcal{H}(G,I)$ on a characteristic function of an $I$ orbit.
This paper is the first in a series of three papers about the module $S(X)^{I}$ over $H(G,I)$. In a future paper we will describe the structure of the module $S(X)^{I}$ in the special case of $X$ being of minimal rank (see [Res10] for a definition). In an additional future paper we will use the results of this paper to answer the question of when the Steinberg representation of $G$ is distinguished with respect to $H$, under the assumption of $H$ being split.
@@ 1.1 Main results
Let $G$ be a connected reductive group split over a $p$-adic field $F$. Assume that the residual characteristic of $F$ is not 2. Let $W_{aff}$ be the affine Weyl group attached to the affine root system of $G$. Let $X=\mathbf{G}(F)/\mathbf{H}(F)$ be a $p$-adic symmetric space.
We construct an action of $W_{aff}$ on the set $I\backslash X$ with finitely many orbits and we use this action to obtain a presentation of the module $S(X)^{I}$ over the Iwahori-Hecke algebra. This generalizes the well known presentation of the Hecke algebra given by Iwahori and Matsomoto. In the group case, where $G=H\times H$ and $H$ is embedded diagonally in $G$, our presentation of $S(H)^{I\times I}=\mathcal{H}(H,I)$ as a $\mathcal{H}(H,I)\times\mathcal{H}(H,I)$ module is exactly the presentation given by [IM65].
Our first main result is the following:
@@
There is a natural action of $W_{aff}$ on $I\backslash X$. This action has finitely many orbits.
We now describe our results in more detail. Recall that a torus $T\subset G$ is called $\sigma$ stable if $\sigma(T)=T$. Let $\mathcal{T}_{\sigma}$ be the set of $\sigma$ stable maximal tori. For each $T\in\mathcal{T}_{\sigma}$ one can attach an affine Weyl group $W_{aff}(T)=N_{G}(T)/T^{0}$. Here, $T^{0}$ is the maximal compact subgroup of $T$. We also attach to $T$ a set $W^{H}_{aff}(T)$ of cosets of $W_{aff}(T)$ with respect to a subgroup $W_{H,aff}(T)=Im(N_{H}(T)\rightarrow N_{G}(T)/T^{0})$
that is naturally embedded into $W_{aff}(T)$. Thus,
$W^{H}_{aff}(T)=W_{aff}(T)/W_{H,aff}(T)$
It is known that $H$ acts on $\mathcal{T}_{\sigma}$ with finitely many orbits (see [HW93]) and we let $A=\mathcal{T}_{\sigma}/H$ be a set of representatives.
Theorem 1.1 from the following description of the set $I\backslash X$.
@@
There is a natural bijection between the $I$ orbits on $X$ and $\cup_{T\in A}W^{H}_{aff}(T)$
@@
The same description was obtained independently by Paul Broussous, see [Bro24].
Using this bijection we define an action of $W_{aff}$, the affine Weyl group of $G$, on the $I$ orbits on $X$. We denote this action by $w\times x$ for $w\in W_{aff}$ and $x\in I\backslash X$ an $I$ orbit on $X$.
We now turn our attention to the description of the action of the Iwahori Hecke algebra $\mathcal{H}(G,I)$ on $S(X)^{I}$. We prove two results regarding this action: the first is a generalization of the classical Iwahori-Matsomoto relations, but it is not a complete description. The second result completes the description but is more involved. We begin with the first result.
The algebra $\mathcal{H}(G,I)$ is generated (as an algebra) by $\{T_{s}|s\in\tilde{\Delta}\}$ and by $\{T_{\omega}|\omega\in\Omega\}$, $\tilde{\Delta}$ being the set of simple reflections inside $W_{aff}$ and $\Omega$ being the group of length zero elements in $W_{aff}$ (see 2.5 for the definition of $\Omega)$. Thus, in order to describe the action of $\mathcal{H}(G,I)$ it is enough to describe the actions of each $T_{s}$ for $s\in\tilde{\Delta}$ and each $T_{\omega}$ for $\omega\in\Omega$.
In section 5 (see Definition 5.1) we define a length function, $l_{\sigma}$, on the set of $I$ orbits on $X$. This length function is similar to the standard length function on $W_{aff}$. We prove the following:
@@
if $l_{\sigma}(s\times x)>l_{\sigma}(x)$ then $T_{s}1_{x}=1_{s\times x}$
if $l_{\sigma}(s\times x)<l_{\sigma}(x)$ then $T_{s}1_{x}=(q-1)1_{x}+q1_{s\times x}$
Unlike in the description of $\mathcal{H}(G,I)$ given by Iwahori and Matsumoto in [IM65], $l_{\sigma}(s\times x)=l_{\sigma}(x)$ is a possibility and in this case the action of $T_{s}$ is more complicated and is not described by the Theorem above. Nevertheless, this result is sufficient to deduce that $S(X)^{I}$ is finitely generated over $\mathcal{H}(G,I)$.
For the second result we relay heavily on the geometry of the Bruhat-Tits Building of $G$. We first prove that $(T_{s}+1)1_{x}$ is constant on its support and we present a description of this support in terms of chambers in the affine building of $G$.
It turns out that there are four options for the size of the support of $(T_{s}+1)1_{x}$. Moreover, we give a way to determine the value of $(T_{s}+1)1_{x}$ on its support by comparing the $l_{\sigma}$ lengths of the elements in the support. To formulate the result we use further notations.
@@
Let $x$ be an $I$ orbit on $X$, $s\in\tilde{\Delta}$ be a simple reflection, $Z\subset I\backslash G/H$ be the support of $(T_{s}+1)1_{x}$.
Denote by $n_{max},n_{min}$ be the number of elements in $Z$ of maximal/minimal $l_{\sigma}$ length respectively. Meaning, let $l_{max}=\text{max}\{l_{\sigma}(z)|z\in Z\}$ and let $n_{max}=\#\{y|l_{\sigma}(y)=l_{max}\}$. $n_{min}$ is defined similarly.
Let $\delta_{x,max}$ be 1 if $l_{\sigma}(x)=l_{max}$ and 00 otherwise. $\delta_{x,min}$ is defined similarly.
@@
$(T_{s}+1)1_{x}$ is constant on its support, denote this constant by $\gamma$. We have $n_{max},n_{min}\in\{1,2\}$. If $l_{\sigma}(x)=l_{\sigma}(s\times x)$ then
@@
In the last section of this paper, we illustrate through the cases of $SL_{2}/T$ and $SL_{2n}/Sp_{2n}$ the concrete results one can obtain from our general theorems.
@@ 1.2 Methods of proof
We begin by noting that the study of $I$ orbits on $X$ can be translated into the study of $H$ orbits on $I\backslash G$. One can identify $I\backslash G$ with the chambers in the affine building of $G$ together with a $\Omega$ coloring. Thus $I\backslash X$ can be replaced with $H$ orbits on the $\Omega$ colored chambers in the affine building $\mathcal{B}_{G}$.
Taking a cue from [DS11], the starting point of our work is the following:
@@
Any chamber in $\mathcal{B}$ is contained in a $\sigma$ stable apartment. This apartment is unique up to conjugation by $H$.
@@
To any element of $x\in I\backslash X$ we attach the characteristic function $1_{x}$, and those form a basis for the vector space $S(X)^{I}$. Here $x$ is thought of as a $\Omega$ colored chamber in $\mathcal{B}_{G}$, the affine building of $G$. A second tool we use, is our length function
$l_{\sigma}:I\backslash X\to\mathbb{N}$
that is defined to be the distance between the chambers $x$ and $\sigma(x)$ in the building $\mathcal{B}_{G}$.
Theorem 1.4 describes the action $T_{s}$ on $1_{x}$, where $s$ is a simple reflection, in the case where $l_{\sigma}(s\cdot x)\neq l_{\sigma}(x)$.
@@
Let $x\in I\backslash G/H,s\in\tilde{\Delta}$ be a simple reflection. Let $\mathcal{C}$ be the chamber fixed by $I$. Choose $g\in G$ that is representative of the orbit $x$ and choose a $\sigma$-stable apartment $\mathcal{A}$ containing both $g\mathcal{C},gs\mathcal{C}$. Finally let $f=g\mathcal{C}\cap gs\mathcal{C}$ and let $\Omega$ be the intersection of all $\sigma$ stable apartments that contain $f$.
@@
  • $l_{\sigma}(s\times x)>l_{\sigma}(x)$ if and only if $g\mathcal{C}\subset\Omega$
  • $l_{\sigma}(s\times x)<l_{\sigma}(x)$ if and only if $gs\mathcal{C}\subset\Omega$
The case $l_{\sigma}(s\times x)=l_{\sigma}(x)$ is more complicated and requires a different approach. We introduce a group $G_{f}$ obtained as the reductive quotient associated to $f\cap\sigma(f)$ (see subsection 2.3 for properties of the reductive quotient). Here $f$ is the face $f=g\mathcal{C}\cap gs\mathcal{C}$. $G_{f}$ is a reductive group of semisimple rank one, we make the observation that the $H$ orbits appearing in the support of $(T_{s}+1)1_{x}$ are in bijection with the orbits of a certain subgroup $\tilde{H}$ of $G_{f}$ on the flag variety of $G_{f}$ (See Proposition 6.6). All this is relevant to the calculation of $T_{s}1_{x}$ because a closer look at the key calculation (see Proposition 4.3):
$(T_{s}+1)1_{x}=\gamma_{f,g}D_{f}$
@@ 1.3 Structure of the paper
In Section 3 we prove Theorem 3.3 describing the set $I\backslash X$ of $I$ orbits on $X$ in terms of quotients of the affine Weyl group. We then introduce an action of $W_{aff}$ on $I\backslash X$ and verify Theorem 1.1.
In Section 4 we preform a key computation describing the action of a Iwahori-Matsumoto generator $T_{s}$ of $\mathcal{H}(G,I)$ on a characteristic function $1_{x}$ for $x\in I\backslash X$. The answer we obtain is in terms of a combinatorial quantity $\gamma$ which is the size of a specific $H$-orbit of chambers in the affine building $\mathcal{B}_{G}$ attached to $x$ and to $s\cdot x$.
@@
In Section 6 we discuses actions of involution on parahoric group schemes and obtain a bijection between $H$ orbits on certain chambers in $\mathcal{B}_{G}$ and the orbits of a symmetric subgroup of $G_{f}$, a reductive quotient, on the flag varieties of $G_{f}$.
@@
In Section 8 we use the results of the previous sections to introduce a generic module $M_{t}$ over the generic Hecke algebra that is described using generator and relations.
@@ 1.4 Related works
Regarding theorem 1.1 we mention that the generalizations of Cartan decomposition to the context of symmetric and spherical spaces was extensively studied. We mention here the work [Off04] of Offen in some special cases as well as the work by [MR09]. In the symmetric spaces case by Delorme and Secherre [DS11] and Benoist and Oh [BO07]. We also mention the work [Sak12] of Sakellaridis considering split groups on a wide class of spherical varieties.
In these works the authors study the orbits of $K_{0}$, a maximal compact subgroup of $G$, on $X$. Particularly in [DS11], Delorme and Secherre, used the affine building of $G$ to give a description of the $K_{0}$ orbits on $X$. Our technique is inspired by their work to give a description of the $I$ orbits on $X$.
In [Kno95] an action of the Weyl group on the Borel orbits of a spherical space is introduced.
The action of $W_{aff}$ we introduce on $I\backslash X$ is an affine version of this in the case of symmetric spaces. Applying our arguments to the spherical building (instead of the affine building) recovers these results in the case of symmetric spaces.
The module $S(X)^{K_{0}}$ over the Hecke algebra $H(G,K_{0})$ was studied extensively, for example by Hironaka [Hir99] for certain spherical and symmetric varieties and by Sakellaridis [Sak13] for a wide class of spherical ones. After determining the spherical functions, they used a Satake type transform in order to obtain a description of the module. Our study of $S(X)^{I}$ for $X$ symmetric is based on geometry and seem to be new. Recently the modules of $I$ fixed vectors of the Gelfand Graev models and other models were considered. See in particular the results of [CS18].
A study of Hecke modules in the algebraic case was initiated in the work of Lusztig-Vogan [LV83] where they study the Grothendieck group of the category of $l$-adic constructible $H$ equivariant sheaves on the flag variety $G/B$ as a module over $H(G.B)$ where $B$ is a Borel subgroup of $G$. Although our methods of proofs are different, our formulas resembles their description.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Let $F$ be a non archimedean local field with residue field $k$. We assume that the characteristic of $k$ is not 2.
[…]
@@
Let $I$ be an Iwahori subgroup of $G$ (defined in [18]). The group $I$ splits the category $\mathcal{M}(G)$ of smooth representations of $G$ (see [4]). The subcategory of representations generated by their $I$ fixed vectors is the so-called principal block. The category of admissible representations generated by their $I$ fixed vectors was studied in [6] and was shown to be equivalent to the category of finite dimensional $\mathcal{H}(G,I)$ modules, the equivalence is given by $V\rightarrow V^{I}$. Here, $\mathcal{H}(G,I)=S(G)^{I\times I}$ is the Hecke Iwahori algebra of $G$ consisting of compactly supported locally constant $I$ bi-invariant functions on $G$.
A first step towards the study of smooth $H$-distinguished representations is the study of $\mathcal{H}(G,I)$ modules with a non-trivial equivariant map from $S(X)^{I}$. Thus, a description of $S(X)^{I}$ as a module over $\mathcal{H}(G,I)$ is a crucial step. In the present work, we do exactly this. Below we describe our main results in detail: Using the affine building $\mathcal{B}$ of $G$ we are able to describe the $I$ orbits on $X$, which give a linear basis of $S(X)^{I}$ over $\mathbb{C}$. Then we describe the action of the Iwahori Matsumoto generators of $\mathcal{H}(G,I)$ on a characteristic function of an $I$ orbit.
This paper is the first in a series of three papers about the module $S(X)^{I}$ over $H(G,I)$. In a future paper, we will describe the structure of the module $S(X)^{I}$ in the special case of $X$ being of minimal rank (see [27] for a definition). In an additional future paper we will use the results of this paper to answer the question of when the Steinberg representation of $G$ is distinguished with respect to $H$, under the assumption of $H$ being split.
@@ 1.1 Main results
Let $W_{aff}$ be the extended affine Weyl group attached to the affine root system of $G$.
[…]
We construct an action of $W_{aff}$ on the set $X/I$, which has finitely many orbits. We use this action to obtain a presentation of the module $S(X)^{I}$ over the Iwahori-Hecke algebra. This generalizes the well known presentation of the Iwahori-Hecke algebra given by Iwahori and Matsumoto. In the group case, where $G=H\times H$ and $H$ is embedded diagonally in $G$, our presentation of $S(H)^{I\times I}=\mathcal{H}(H,I)$ as a $\mathcal{H}(H,I)\times\mathcal{H}(H,I)$ module is exactly the presentation given in [18].
Our first main result is the following, it is proven in Subsection 3.2.
@@
There is an explicit action of $W_{aff}$ on $X/I$. This action has finitely many orbits.
We now describe our results in more detail. A torus $T\subset G$ is called $\sigma$ stable if $\sigma(T)=T$. Let $\mathcal{T}_{\sigma}$ be the set of $\sigma$ stable maximal tori. For each $T\in\mathcal{T}_{\sigma}$ one can attach an extended affine Weyl group $W_{aff}(T)=N_{G}(T)/T^{0}$. Here, $T^{0}$ is the unique maximal compact subgroup of $T$. We also attach to $T$ a set $W^{H}_{aff}(T)$ of cosets of $W_{aff}(T)$ with respect to a subgroup $W_{H,aff}(T)=Im(N_{H}(T)\rightarrow N_{G}(T)/T^{0})$ that is naturally embedded into $W_{aff}(T)$.
Thus,
[…]
It is known that $H$ acts on $\mathcal{T}_{\sigma}$ with finitely many orbits (see [15]) and we let $A=H\backslash\mathcal{T}_{\sigma}$ be a set of representatives.
Theorem 1.1 follows from the following description of the set $X/I$.
@@
There is an explicit bijection between the $I$ orbits on $X$ and $\cup_{T\in A}W^{H}_{aff}(T)$
The proof is given in Section 3 in Theorem 3.4.
@@
The same description was obtained independently by Paul Broussous, see [7].
Using this bijection we define a right action of $W_{aff}$, the extended affine Weyl group of $G$, on the $I$ orbits on $X$.
We denote this action by $x\times w$ for $w\in W_{aff}$ and $x\in X/I$.
We now turn our attention to the description of the action of the Iwahori Hecke algebra $\mathcal{H}(G,I)$ on $S(X)^{I}$. We prove two results regarding this action: the first is a generalization of the classical Iwahori-Matsumoto relations, but it is not a complete description. The second result completes the description but is more involved. We begin with the first result.
Let $\tilde{\Delta}$ be the set of simple reflections inside $W_{aff}$ and let $\Omega$ be the group of length zero elements in $W_{aff}$ (see 2.5 for the definition of $\Omega)$.
For any $w\in W_{aff}$ we denote $T_{w}=1_{IwI}\in\mathcal{H}(G,I)$ the characteristic function of $IwI$.
The algebra $\mathcal{H}(G,I)$ is generated (as an algebra) by $\{T_{s}|s\in\tilde{\Delta}\}$ and by $\{T_{o}|o\in\Omega\}$. Thus, in order to describe the action of $\mathcal{H}(G,I)$ it is enough to describe the actions of each $T_{s}$ for $s\in\tilde{\Delta}$ and each $T_{o}$ for $o\in\Omega$.
In Section 5 (see Definition 5.1) we define a length function, $l_{\sigma}$, on the set of $I$ orbits on $X$. This length function is similar to the standard length function on $W_{aff}$. We prove the following (see Theorem 5.3).
@@
if $l_{\sigma}(x\times s)>l_{\sigma}(x)$ then $1_{x}T_{s}=1_{x\times s}$
if $l_{\sigma}(x\times s)<l_{\sigma}(x)$ then $1_{x}T_{s}=(q-1)1_{x}+q1_{x\times s}$
Unlike in the description of $\mathcal{H}(G,I)$ given by Iwahori and Matsumoto in [18], $l_{\sigma}(x\times s)=l_{\sigma}(x)$ is a possibility and in this case the action of $T_{s}$ is more complicated and is not described by the Theorem above. Nevertheless, this result is sufficient to deduce that $S(X)^{I}$ is finitely generated over $\mathcal{H}(G,I)$.
For the second result we rely heavily on the geometry of the affine building of $G$. We first prove that $1_{x}(T_{s}+1)$ is constant on its support and we present a description of this support in terms of chambers in the affine building of $G$.
It turns out that there are four options for the size of the support of $1_{x}(T_{s}+1)$. Moreover, we give a way to determine the value of $1_{x}(T_{s}+1)$ on its support by comparing the $l_{\sigma}$ lengths of the elements in the support. To formulate the result we need further notations.
@@
Let $x$ be an $I$ orbit on $X$, $s\in\tilde{\Delta}$ be a simple reflection, and let $Z\subset H\backslash G/I$ be the support of $1_{x}(T_{s}+1)$.
Denote by $n_{max},n_{min}$ be the number of elements in $Z$ of maximal/minimal $l_{\sigma}$ length respectively. Meaning, let $l_{max}=\text{max}\{l_{\sigma}(z)|z\in Z\}$ and let $n_{max}=\#\{y\in Z|l_{\sigma}(y)=l_{max}\}$. The value $n_{min}$ is defined similarly.
Let $\delta_{x,max}$ be 1 if $l_{\sigma}(x)=l_{max}$ and 0 otherwise. The value $\delta_{x,min}$ is defined similarly.
@@
$1_{x}(T_{s}+1)$ is constant on its support, denote this constant by $\gamma$. We have $n_{max},n_{min}\in\{1,2\}$. If $l_{\sigma}(x)=l_{\sigma}(x\times s)$ then
@@
[…]
@@ 1.2 Methods of proof
We begin by noting that the study of $I$ orbits on $X$ can be translated into the study of $H$ orbits on $G/I$. One can identify $G/I$ with the set of chambers in the affine building of $G$ together with a $\Omega$ coloring. Thus $X/I$ can be replaced with $H$ orbits on the $\Omega$ colored chambers in the affine building $\mathcal{B}_{G}$.
Taking a cue from [12], the starting point of our work is the following:
@@
Any chamber in $\mathcal{B}$ is contained in a $\sigma$ stable apartment. This apartment is unique up to an action by an element of $H$.
@@
To any element of $x\in X/I$ we attach the characteristic function $1_{x}$. These functions form a basis for the vector space $S(X)^{I}$. Here $x$ is thought of as a $H$ orbit of $\Omega$ colored chambers in $\mathcal{B}_{G}$, the affine building of $G$. A second tool we use, is our length function
$l_{\sigma}:X/I\to\mathbb{N}$
that is defined to be the distance between a representative $\mathcal{C}_{x}$ of the orbit $x$, and $\sigma(\mathcal{C}_{x})$ in the building $\mathcal{B}_{G}$.
Theorem 1.4 describes the action $T_{s}$ on $1_{x}$, where $s$ is a simple reflection, in the case where $l_{\sigma}(x\times s)\neq l_{\sigma}(x)$.
@@
Let $x\in H\backslash G/I$ and let $s\in\tilde{\Delta}$ be a simple reflection. Let $\mathcal{C}$ be the chamber fixed by $I$. Choose $g\in G$ a representative of the orbit $x$ and choose a $\sigma$-stable apartment $\mathcal{A}$ containing both $g\mathcal{C},gs\mathcal{C}$. Finally let $f=g\mathcal{C}\cap gs\mathcal{C}$ and let $\Pi$ be the intersection of all $\sigma$ stable apartments that contain $f$.
@@
  • $l_{\sigma}(x\times s)>l_{\sigma}(x)$ if and only if $g\mathcal{C}\subset\Pi$.
  • $l_{\sigma}(x\times s)<l_{\sigma}(x)$ if and only if $gs\mathcal{C}\subset\Pi$.
The case $l_{\sigma}(x\times s)=l_{\sigma}(x)$ is more complicated and requires a different approach. We introduce a group $G_{f}$ obtained as the reductive quotient associated to $f\cap\sigma(f)$ (see subsection 2.3 for properties of the reductive quotient). The group $G_{f}$ is a reductive group of semisimple rank one, we make the observation that the $H$ orbits appearing in the support of $1_{x}(T_{s}+1)$ are in bijection with the orbits of a certain subgroup $\tilde{H}$ of $G_{f}$ on the flag variety of $G_{f}$ (See Proposition 6.6). All this is relevant to the calculation of $1_{x}T_{s}$ because a closer look at the key calculation (see Proposition 4.3):
$1_{x}(T_{s}+1)=\gamma_{f,g}D_{f}$
@@ 1.3 Structure of the paper
In Section 3 we prove Theorem 3.4 describing the set $X/I$ of $I$ orbits on $X$ in terms of quotients of the extended affine Weyl group. We then introduce an action of $W_{aff}$ on $X/I$ and verify Theorem 1.1.
In Section 4 we perform a key computation describing the action of a Iwahori-Matsumoto generator $T_{s}$ of $\mathcal{H}(G,I)$ on a characteristic function $1_{x}$ for $x\in X/I$. The answer we obtain is in terms of a combinatorial quantity $\gamma$ which is the size of a specific $H$-orbit of chambers in the affine building $\mathcal{B}_{G}$ attached to $x$ and to $x\times s$.
@@
In Section 6 we discuss actions of involution on parahoric group schemes and obtain a bijection between $H$ orbits on certain chambers in $\mathcal{B}_{G}$ and the orbits of a symmetric subgroup of $G_{f}$, a reductive quotient, on the flag varieties of $G_{f}$.
@@
In Section 8 we use the results of the previous sections to introduce a generic module $M_{t}$ over the generic Iwahori-Hecke algebra that is described using generators and relations.
@@ 1.4 Related works
Regarding Theorem 1.1, we mention that the generalizations of the Cartan decomposition to the context of symmetric and spherical spaces was extensively studied. We mention here the work [26] of Offen in some special cases as well as the work by [23]. In the symmetric spaces case by Delorme and Secherre [12] and Benoist and Oh [2]. We also mention the work [29] of Sakellaridis considering split groups on a wide class of spherical varieties.
In these works the authors study the orbits of $K_{0}$, a maximal compact subgroup of $G$, on $X$. Particularly in [12], Delorme and Secherre, used the affine building of $G$ to give a description of the $K_{0}$ orbits on $X$. Our technique is inspired by their work to give a description of the $I$ orbits on $X$.
In [28] an action of the Weyl group on the Borel orbits of a symmetric space is introduced. In [20] this constructions was extended to the action of the Weyl group on the Borel orbits of any spherical variety.
The action of $W_{aff}$ we introduce on $X/I$ is an affine version of this in the case of symmetric spaces. Applying our arguments to the spherical building (instead of the affine building) recovers these results in the case of symmetric spaces.
The module $S(X)^{K_{0}}$ over the Hecke algebra $H(G,K_{0})$ was studied extensively, for example by Hironaka [17] for certain spherical and symmetric varieties and by Sakellaridis [30] for a wide class of spherical ones. After determining the spherical functions, they used a Satake type transform in order to obtain a description of the module. Our study of $S(X)^{I}$ for $X$ symmetric is based on geometry and seems to be new. Recently the modules of $I$ fixed vectors of the Gelfand Graev models and other models were considered. See in particular the results of [10].
A study of Hecke modules in the algebraic case was initiated in the work of Lusztig-Vogan [22] where they study the Grothendieck group of the category of $l$-adic constructible $H$ equivariant sheaves on the flag variety $G/B$ as a module over $H(G,B)$ where $B$ is a Borel subgroup of $G$. A version of this for the Iwahori case was obtained in [11]. Although our methods of proof are different, our formulas resemble their description.